OpenStudy (jadedry):

Trig Equations + General Solutions. (Almost complete, need help with last few lines)

OpenStudy (jadedry):

Solve the following equations and give angles between 0 and 360. \[\sin 2\theta -1 = \cos 2 \theta\] So far I've completed the following! \[2 \sin \theta \cos \theta = 2 \cos^2 \theta\] =\[= \frac{ 2 \sin \theta \cos \theta }{ 2 \cos ^2 \theta } = 1 = \frac{ 2 \sin \theta }{ 2 \cos \theta} = \tan \] Since tan = 1, tan = 45 degrees, so the general solution is 180n + 45 However, my book says this is only /half/ the answer. The other solution is 180n + 90, how so? Thanks in advance!

OpenStudy (irishboy123):

\(2 \sin \theta \cos \theta = 2 \cos^2 \theta\) can be true for \(cos \theta = 0\) too

zepdrix (zepdrix):

In general, try to avoid "cancelling things out". You lose information when you do that. Rely more heavily on factoring and other techniques. \[\large\rm \sin \theta \cos \theta=\cos^2\theta\]Subtraction gives us\[\large\rm 0=\cos^2\theta-\sin \theta \cos \theta\]Pulling a cosine out of each term,\[\large\rm 0=\cos \theta(\cos \theta-\sin \theta)\]

zepdrix (zepdrix):

And then apply your Zero-Factor Property,\[\large\rm 0=\cos \theta,\qquad\qquad\qquad 0=(\cos \theta - \sin \theta)\]

OpenStudy (jadedry):

@Zepdrix @Irishboy123 Sorry guys! I didn't see your replies earlier. I see what you mean, thanks you both for your help! c: